OpenStudy (anonymous):

Julie wants to have a 90 average in her math class at the end of the year. She is trying to determine what she needs to make on her final exam for this to work. Everything is weighted equally and she currently has a 85 for her quiz average, a 80 for her homework average, a 92 for her test average, and a 98 for her project average. What does she need to make on her final exam for her to end up with a 90 in the class?

OpenStudy (jacobbirddog):

she needs to be a flipping genius.

OpenStudy (anonymous):

oh :3

OpenStudy (anonymous):

that's not one of my answer choices..

OpenStudy (campbell_st):

well its simply a case of summing the know marks... and knowing there average has to be 90... so \[90 = \frac{ 80 + 85 + 92 + 98 + x}{5}\] solve for x, where x is the exam mark

OpenStudy (texaschic101):

you type too fast...

OpenStudy (texaschic101):

lol

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

:(

OpenStudy (campbell_st):

ok.... so you know here scores for other tasks add to 355... so you need to find the average of the 5 types of tasks including the exam... and they need to average 90 so then \[90 = \frac{355 + x}{5}\] multiply both sides of the equation by 5 450 = 355 + x

OpenStudy (campbell_st):

the student needs the sum of all her assessment tasks to be 450, so she can average 90... so you need to solve the equation 450 = 355 + x x will be what's needed in the final exam.

OpenStudy (anonymous):

YES!! <3 thank you

OpenStudy (texaschic101):

To find the average of something, you have to add up the items, then divide by how many items there are... (80 + 85 + 92 + 98 + x) / 5 = 90 (there is 5 grades, we just don't know 1 of them) (355 + x) / 5 = 90 --- multiply by 5 355 + x = 450 take it from here and solve for x, the final grade